# Proficient Graph Neural Network Design by Accumulating Knowledge on Large Language Models

### v2

@@ 3.1 Connect Graph Topology to GNN Performance
Performance-oriented modeling of discrete graph data is essential to quickly capture data priors from massive benchmarks to construct trustworthy meta-knowledge, a key challenge for LLMs due to their inability to naturally interpret graph-structured data [16]. As discussed in Section 2.2, existing LLM-based methods rely heavily on user-provided task descriptions, which fail to capture the key graph topology (e.g., heterophily) that frequently inspires specific GNN designs in the vast graph learning literature [54, 47]. Our studies in Figure 1(b),  3(b) and Appendix D.2 demonstrate that semantic or purely descriptive inputs (e.g., is a citation graph) frequently mislead LLMs, causing “artificial hallucination” and ineffective solutions. Such descriptive inputs introduce significant external noise, undermining the reliability of generic LLM suggestions when connecting graphs to model performance. Thus, it is necessary to establish an empirical foundation that captures fine-grained graph-GNN-performance mappings, enabling the construction of trustworthy meta-knowledge for data-aware model design.
Table 1: 16 graph topological properties
Clustering Betweenness Density Degree centrality
Closeness Degree Edge count Graph diameter
Shortest path Assortativity Eigenvector Dimensionality
Node count Diversity Connected Label homophily
@@
To empirically study the missing meta-level understanding about graph-GNN-performance, we start with 16 graph properties $\{g_{k}\}_{k=1}^{16}$ from the literature (see Table 1 and Appendix A.2) to infer graph similarity from two aspects: 1) the statistical distance ranking $\mathcal{SR}(i,k)$ based on graph properties $g_{k}$, and 2) the empirical performance ranking $\mathcal{ER}(i)$ of transferred top-performing GNNs from other graphs $\{G^{j}\mid j\neq i\}$ to anchor graph $G^{i}$. The concrete definition is presented in Appendix A.2.2. To assess how reliably specific graph properties predict successful GNN design knowledge transfers, we introduce the concept of graph property confidence for $g_{k}$ on $G^{i}$ through Kendall’s $\tau$ Coefficient [23]:
$I(G^{i},g_{k}):=\text{KendallCorr}\big{(}\mathcal{SR}(i,k),\mathcal{ER}(i)\big{)},$ (2)
@@
where $I(G^{i},g_{k})$ quantifies the correlation between the topological distance measured by $g_{k}$ and the empirical performance differences in transferring different top-performing GNNs from $\{G^{j}\mid j\neq i\}$ to $G^{i}$. A higher $I(G^{i},g_{k})$ indicates that $g_{k}$ is a stronger indicator of effective model design transfers, thus playing a critical role in meta-learning with LLMs to design data-aware models for unseen tasks.
@@
Empirical Filtering as Trustworthy Foundation. Recognizing those empirical insights on external noise and inherent property variability, we first utilize empirical filtering to preliminarily identify reliably influential graph properties by averaging their predictive confidences $I(G^{i},g_{k})$ across benchmark graphs as $\bar{I}(g_{k})=\frac{1}{n}\sum_{i=1}^{n}I(G^{i},g_{k}).$ This process ensures that our property set built from the literature is also empirically grounded, reducing susceptibility to external noise and inherent property variability. Then, based on the empirical observation in Figure 6(b), we develop a filter $\mathcal{F}(G^{i},\bar{I})$ to correlate the modeling of discrete graph data with transferred model performance, selecting the Top-$N_{f}$ properties with the highest $\bar{I}(g_{k})$. This leads to the alternative formulation of Equation 1:
$\mathbb{P}(\theta_{i}|\mathcal{F}(G^{u}))=\frac{\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i})\,\mathbb{P}(\theta_{i})}{\mathbb{P}(\mathcal{F}(G^{u}))}.$ (3)
This mechanism mitigates the typical LLM failure of “artificial hallucinations” often triggered by unprofessional graph modeling with negligible computational cost offline. To accumulate new knowledge, we design a self-evaluation bank $BE$, which stores measured graph properties and similarity rankings for new graphs on the fly, allowing $\mathcal{F}$ to scalably evolve over time by integrating new empirical evidence. Grounded in empirical transferability, this mechanism serves as a trustworthy foundation for capturing graph-to-graph ($G^{i}$ to $G^{j}$) meta-priors that underpin the correlations between graphs and high-performing GNN models, but still calling for an effective property variability handler.
3.2 Knowledge Retrieval Module
Proficiency in GNN design largely depends on capturing the meta-knowledge in the form of the posterior belief for unseen tasks. However, even after empirically filtering and establishing a general correlation between graph properties and model performance, solving pseudo-likelihood $\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i})$ remains a resource-intensive accumulation of testing various $\theta_{i}$ on $G^{u}$. To overcome it, we novelly propose to utilize the known likelihood $\mathbb{P}(\mathcal{F}(G^{i})|\theta_{i})$ on the benchmark datasets to form a practical proxy for $\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i})$ on the unseen. Eliciting such fine-grained model design knowledge paves the way for efficient approximation of the posterior $\mathbb{P}(\theta_{i}|\mathcal{F}(G^{u}))$ with accumulated prior knowledge. Finding such a practical proxy for the unseen task among the massive benchmark knowledge requires an unsupervised task similarity formulation that properly assesses the likelihood of successful model design knowledge transfer from $G^{i}$ to $G^{u}$, which has to reflect the performance outcomes $\mathcal{H}(\cdot,\cdot;G^{u})$.
Definition 1 (Task Similarity for Model Design Knowledge Transfer)
Given an unseen graph $G^{u}$ and two benchmark graphs $G^{i}$ and $G^{j}$, each associated with $N_{m}$ top-performing models $\{\theta^{*}_{m}\}_{m=1}^{N_{m}}\in\Theta$, an effective task similarity $\mathcal{S}(\cdot,\cdot)$ for transfering model design knowledge satisfies:
$\mathcal{S}(G^{u},G^{i})\geq\mathcal{S}(G^{u},G^{j})$ (4)
$\Rightarrow$ [($\mathbb{P}\Bigl{[}\mathcal{H}(\{\theta^{*}_{im}\},\cdot,G^{u})\geq\mathcal{H}(\{\theta^{*}_{jm}\},\cdot,G^{u})\Bigr{]}\geq\delta.$
where ($\mathcal{H}(\{\theta^{*}_{im}\},\cdot,G^{u})$ denotes the performance of the transferred top-performing pattern $\{\theta^{*}_{im}\}$ from benchmark $G^{i}$ on the unseen $G^{u}$, and $\delta$ is a confidence threshold. This ensures that higher task similarity between $G^{u}$ and $G^{i}$ implies a larger likelihood of high-performing transfer from $G^{i}$ to $G^{u}$.
To design such performance-oriented task similarity between $G^{u}$ and $G^{i}$ unsupervisedly, we first apply the filter $\mathcal{F}(\cdot)$ (initialized with accumulated benchmark knowledge in Section 3.1) on $G^{u}$ and $G^{i}$ to establish an empirical foundation based on performance-oriented confidence $\bar{I}(g_{k})$. However, as discussed in Section 3.1, graph properties inherently exhibit varying levels of confidences across graphs, making it challenging to transfer knowledge reliably with standalone statistical properties. To account for this inherent variability, we further introduce the adaptive importance $w^{u}_{k}$ for each property $g_{k}$ specifically tailored to the unseen graph $G^{u}$ and formulate our unsupervised task similarity as:
$\mathcal{S}(G^{u},G^{i})=\frac{1}{N_{f}}\sum_{k=1}^{N_{f}}w^{u}_{k}\cdot\frac{\bar{I}(g_{k})}{1+\hat{d_{k}^{ui}}},$ (5)
where $N_{f}$ is the number of influential graph properties selected via empirical filtering and $\bar{I}(g_{k})/(1+\hat{d_{k}^{ui}})$ measures the confident performance similarity of property $g_{k}$ between graphs $G^{u}$ and $G^{i}$.
@@ 3.2 Knowledge Retrieval Module
Adaptive Knowledge Elicitation and Alignment. While $\bar{I}(g_{k})$ is grounded in the observations from benchmarks, their adaptation to unseen tasks requires finding the adaptive importance of property $w^{u}_{k}$, which cannot be observed without testing on $G^{u}$. Notably, such adaptive importance of property for a given graph commonly motivates specific GNN designs in the graph learning literature [54, 47], e.g., GraphSAGE [18] handles high-degree graphs. Therefore, we novelly propose to leverage LLMs—pre-trained on a vast corpus of graph learning literature—as meta-controllers to adaptively elicit $w^{u}_{k}$ for $G^{u}$ by aligning the task descriptions, filtered graph properties, and high-performing architecture patterns in textual space, followed by contextual reasoning [31]. This process ensures that the data-specific weights $w^{u}_{k}$ reflect the interpretable literature insights between graph properties, GNNs, and performance, aligning with general empirical evidence to construct trustworthy meta-knowledge that empowers LLMs to quickly design effective, data-aware models for unseen tasks.
@@
Knowledge Pool Construction. Using the computed task similarity set $\mathcal{S}^{u}$ of $G^{u}$, we rank the benchmark graphs (Appendix A.3) and select the Top-$N_{s}$ most similar sources. The knowledge pool $\mathcal{K}$ for $G^{u}$ is then constructed by aggregating the top-performing models from these selected sources:
$\mathcal{K}=\bigcup_{G^{i}\in\text{Top-}{N_{s}}(\mathcal{S}^{u})}\{(G^{i},\{\theta_{im}^{*}\}_{m=1}^{N_{m}})\},$ (6)
@@
The top-performing models $\{\theta_{i}^{*}\}$ act as a truncated prior $\mathbb{P}(\theta_{i}^{*})$, summarizing accumulated design knowledge from benchmark $G^{i}$, while task similarity $\mathcal{S}(G^{u},G^{i})$ serves as a gating module for solving pseudo-likelihood $\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i}^{*})$, indicating the effectiveness of $\theta_{i}^{*}$ to the unseen graph $G^{u}$.
3.3 Model Suggestion and Refinement
Overall, our formulation of Knowledge Pool $\mathcal{K}$ represents the approximated posterior $\mathbb{P}(\theta_{i}^{*}|\mathcal{F}(G^{u}))$ that prioritizes models that align with data-aware evidence. This Bayesian-inspired process ensures solid usage of constructed meta-knowledge for designing models in both one-shot and a few attempts. We present our detailed procedure in Algorithm 1 and prompt designs in code implementation.
Initial Model Suggestion. By supplying contextual meta-knowledge of high similarities to the unseen graph $G^{u}$ from $\mathcal{K}$, we leverage LLM as a surrogate meta-controller to samples $N_{s}$ initial proposals $\{\theta_{ui}\}_{i=1}^{N_{s}}$ based on the approximated posterior: $\theta_{ui}\leftarrow\mathcal{LLM}_{IMS}(\mathcal{F}(G^{u},\bar{I}),\mathcal{K}_{i}),$ where $\mathcal{K}_{i}$ represents the specific knowledge from the $i$-th similar benchmark graph in $\mathcal{K}$. Compared to existing AutoGNNs, our method avoids any model training on the unseen before giving promising designs, combining fine-grained literature and empirical insights for immediate model delivery.
Knowledge-Driven Refinement. To further optimize initial model proposals, we design a structured, knowledge-driven refinement strategy inspired by expert-like exploration and exploitation tradeoffs, which is illustrated in Appendix A.4. Our local search strategy iteratively improves models using the constructed meta-knowledge of high similarities to the unseen from $\mathcal{K}$. This refinement process balances exploration (testing diverse configurations) and exploitation (focusing on high-potential candidates), leveraging prior knowledge in textual space to optimize model designs efficiently.
4 Experiments
4.1 Experimental Settings
Task and Datasets. We conduct experiments on the node classification using 8 benchmark datasets from NAS-Bench-Graph [35]: Cora, Citeseer, PubMed [37], CS, Physics, Photo, Computer [38], and ogbn-arXiv [20]. Besides, we include 3 extra datasets for out-of-distribution (e.g., heterophily) generalization: DBLP [3], Flickr [56], and Actor [34]. The statistics are summarized in Appendix B.1.
Baselines. We compare DesiGNN with 4 categories of 17 baselines: (1) Manually Designed GNNs: 7 GNNs [24, 42, 18, 52, 10, 2, 32] showcasing human experts’ efforts in designing tailored models for specific types of graphs; (2) Classical Automated Approaches: 3 typical automated search strategies [27, 36, 65], plus 2 advanced AutoGNN systems [15, 62], all operating within the same search space as DesiGNN for fairness; (3) Meta-based Approaches: 3 meta-based, semi-supervised methods [54, 35, 5] that recommend models from the most similar benchmark graphs or based on meta-knowledge after testing anchor models on the unseen task; (4) LLM-Based Approaches: 2 recent SOTA LLM-Based methods [43, 12] that leverage LLMs for iterative architecture search.
Evaluation Metrics. To ensure reliability, we report the average accuracy and standard deviation over 10 runs. For automated methods, we also analyze the best-so-far accuracy after validating 1-30 model proposals to measure short-run efficiency. To assess efficiency in the case of varying LLM API traffic, we set the basic time unit as the time to validate a model proposal on the unseen dataset.
Implementation Details. The DesiGNN code is implemented in PyTorch and LangChain (GPT-4). For clarity, Desi-Init refers to initial model proposals generated in Section 3.3, while DesiGNN denotes the full pipeline, including model refinement in Section 3.3. Our experiments use NAS-Bench-Graph [35] for accumulated knowledge and a unified search space. To prevent data leakage, when a benchmark dataset is treated as unseen, it is completely excluded from the knowledge retrieval process and anonymized, ensuring fair evaluation. Out-of-distribution datasets (DBLP, Flickr, Actor) have access to all benchmark graphs as knowledge sources. We have validated that LLMs, given only dataset descriptions, lack prior knowledge of effective architectures in the NAS-Bench-Graph search space in Appendix D.1. Hyperparameters and more details are in Appendix B.1.
4.2 Main Results
@@
Table 2: Initial performance comparison (accuracy) of GNNs designed by different methods. $N_{s}=3$ for Desi-Init. Best overall results are in bold, and best among each category are underlined.

### v3

@@ 3.1. Connect Graph Topology to GNN Design
Performance-oriented modeling of discrete graph data is essential to quickly capture data priors from massive benchmarks to construct trustworthy meta-knowledge, a key challenge for LLMs due to their inability to naturally interpret graph-structured data (Guo et al., 2023). As discussed in Section 2.2, existing LLM-based methods rely heavily on user-provided task descriptions, which fail to capture key graph topologies (e.g., heterophily) that often motivate specific GNN designs in the vast graph learning literature (You et al., 2020; Wang et al., 2023b). Our studies in Figures 1(b),  3(b), and Section 4.3 demonstrate that semantic or purely descriptive inputs (e.g., a citation graph) frequently mislead LLMs, leading to “artificial hallucination” and ineffective solutions. Such descriptive inputs introduce significant external noise, undermining the reliability of generic LLM suggestions when connecting graphs to model performance. Thus, it is necessary to establish an empirical foundation that captures fine-grained graph-GNN-performance mappings, enabling the construction of trustworthy meta-knowledge for data-aware model design.
@@
To empirically study the missing meta-level understanding about graph-GNN-performance, we start with 16 graph properties $\{g_{k}\}_{k=1}^{16}$ from the literature: clustering, betweenness, density, degree centrality, closeness, degree, edge count, graph diameter, shortest path, assortativity, eigenvector, dimensionality, node count, diversity, connected component, and label homophily. We infer graph similarity from two aspects: 1) the statistical distance ranking $\mathcal{SR}(i,k)$ based on graph properties $g_{k}$, and 2) the empirical performance ranking $\mathcal{ER}(i)$ of transferred top-performing GNNs from other graphs $\{G^{j}\mid j\neq i\}$ to anchor graph $G^{i}$. To assess how reliably specific graph properties predict successful GNN design knowledge transfers, we introduce the concept of graph property confidence for $g_{k}$ on $G^{i}$ through Kendall’s $\tau$ Coefficient (Kendall, 1938):
(2) $I(G^{i},g_{k}):=\text{KendallCorr}\big(\mathcal{SR}(i,k),\mathcal{ER}(i)\big),$
@@
where $I(G^{i},g_{k})$ quantifies the correlation between the topological distance measured by $g_{k}$ and the empirical performance differences in transferring different top-performing GNNs from $\{G^{j}\mid j\neq i\}$ to $G^{i}$. A higher $I(G^{i},g_{k})$ indicates that $g_{k}$ is a stronger indicator of effective model design transfers, thus playing a key role in meta-learning with LLMs to design data-aware models for unseen tasks.
Variability in Property Confidence. To validate the reliability of graph properties for GNN performance, we assess 16 confidence scores $\{I(G^{i},g_{k})\}_{k=1}^{16}$ for each anchor graph $G^{i}$ against the other 8 graphs in the benchmark (Qin et al., 2022). Figure 3(a) shows that confidence $I(G^{i},g_{k})$ fluctuates widely across diverse graphs. e.g., while $I(\text{Physics},g_{16}\text{=homophily})$ is positive, the same property exhibits negative correlations for other graphs. This inherent variability highlights that specific graph properties are strong indicators of transferability for certain graphs but unreliable for others. Notably, this directly exposes why generic LLM suggestions fail to capture fine-grained insights—LLMs cannot implicitly discern when particular properties matter without explicit, empirically validated priors.
Empirical Filtering as Trustworthy Foundation. Recognizing those empirical insights on external noise and inherent property variability, we first use empirical filtering to preliminarily identify graph properties that are reliably influential by averaging their predictive confidences $I(G^{i},g_{k})$ across benchmark graphs as $\bar{I}(g_{k})=\frac{1}{n}\sum_{i=1}^{n}I(G^{i},g_{k}).$ This process ensures that our property set built from the literature is also empirically grounded, reducing susceptibility to external noise and inherent property variability. Then, based on the empirical study in Figure 4, we develop a filter $\mathcal{F}(G^{i},\bar{I})$ to correlate the discrete graph modeling with transferred model performance, selecting the Top-$N_{f}$ properties with the highest $\bar{I}(g_{k})$. This leads to the alternative formulation of Equation 1:
(3) $\mathbb{P}(\theta_{i}|\mathcal{F}(G^{u}))=\frac{\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i})\,\mathbb{P}(\theta_{i})}{\mathbb{P}(\mathcal{F}(G^{u}))}.$
This mechanism mitigates the typical LLM failure of “artificial hallucinations”—often triggered by unprofessional graph modeling—at negligible offline computational cost. To accumulate new knowledge, we design a self-evaluation bank $BE$ that stores measured graph properties and similarity rankings for new graphs on the fly, enabling $\mathcal{F}$ to scalably evolve over time by integrating new empirical evidence. Grounded in empirical transferability, this mechanism serves as a trustworthy foundation for capturing graph-to-graph ($G^{i}$ to $G^{j}$) meta-priors that underpin the correlations between graphs and high-performing GNN models, yet it still calls for an effective property variability handler.
Figure 4. Average confidence of graph properties.
3.2. Knowledge Retrieval Module
Proficiency in GNN design largely depends on capturing meta-knowledge in the form of the posterior belief for unseen tasks. However, even after empirically filtering and establishing a general correlation between graph properties and model performance, solving pseudo-likelihood $\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i})$ remains a resource-intensive accumulation of testing various $\theta_{i}$ on $G^{u}$. To overcome it, we novelly propose to leverage the known likelihood $\mathbb{P}(\mathcal{F}(G^{i})|\theta_{i})$ on benchmark datasets to form a practical proxy for $\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i})$ on the unseen. Eliciting such fine-grained model design knowledge paves the way for an efficient approximation of the posterior $\mathbb{P}(\theta_{i}|\mathcal{F}(G^{u}))$ with accumulated prior knowledge. Finding such a proxy for the unseen task among the massive benchmark knowledge requires an unsupervised task similarity formulation that properly assesses the likelihood of successful model design knowledge transfer from $G^{i}$ to $G^{u}$, which has to reflect the performance outcomes $\mathcal{H}(\cdot,\cdot;G^{u})$.
Definition 3.1 (Task Similarity for Model Design Knowledge Transfer).
Given an unseen graph $G^{u}$ and two benchmark graphs $G^{i}$ and $G^{j}$, each associated with $N_{m}$ top-performing models $\{\theta^{*}_{m}\}_{m=1}^{N_{m}}\in\Theta$, an effective task similarity $\mathcal{S}(\cdot,\cdot)$ for transferring model design knowledge satisfies:
(4) $\mathcal{S}(G^{u},G^{i})\geq\mathcal{S}(G^{u},G^{j})$
$\Rightarrow$ $\mathbb{P}\Bigl[\mathcal{H}(\{\theta^{*}_{im}\},\cdot,G^{u})\geq\mathcal{H}(\{\theta^{*}_{jm}\},\cdot,G^{u})\Bigr]\geq\delta.$
where $\mathcal{H}(\{\theta^{*}_{im}\},\cdot,G^{u})$ denotes the performance of the transferred top-performing pattern $\{\theta^{*}_{im}\}$ from benchmark $G^{i}$ on the unseen $G^{u}$, and $\delta$ is a confidence threshold. This ensures that higher task similarity between $G^{u}$ and $G^{i}$ implies a larger likelihood of high-performing transfer from $G^{i}$ to $G^{u}$.
To design unsupervised task similarity between $G^{u}$ and $G^{i}$, we first apply the filter $\mathcal{F}(\cdot)$ on $G^{u}$ and $G^{i}$ to establish an empirical foundation based on performance-oriented confidence $\bar{I}(g_{k})$. However, as discussed in Section 3.1, graph properties inherently exhibit varying levels of confidence across graphs, making it challenging to transfer knowledge reliably with standalone statistical properties. To account for inherent variability, we further introduce the adaptive importance $w^{u}_{k}$ for each property $g_{k}$ specifically tailored to the unseen graph $G^{u}$ and define our unsupervised task similarity as:
(5) $\mathcal{S}(G^{u},G^{i})=\frac{1}{N_{f}}\sum_{k=1}^{N_{f}}w^{u}_{k}\cdot\frac{\bar{I}(g_{k})}{1+\hat{d_{k}^{ui}}}$,
@@ 3.2. Knowledge Retrieval Module
Adaptive Knowledge Elicitation and Alignment. While $\bar{I}(g_{k})$ is grounded in the observations from benchmarks, their adaptation to unseen tasks requires finding the adaptive importance of property $w^{u}_{k}$, which cannot be observed without testing on $G^{u}$. Notably, such adaptive importance of property for a given graph commonly motivates specific GNN designs in the graph learning literature (You et al., 2020; Wang et al., 2023b), e.g., GraphSAGE (Hamilton et al., 2017) handles high-degree graphs. Therefore, we novelly propose to leverage LLMs—pre-trained on a vast corpus of graph learning literature—as meta-controllers to adaptively elicit $w^{u}_{k}$ for $G^{u}$ by aligning task descriptions, filtered graph properties, and high-performing architecture patterns in textual space, followed by contextual reasoning (Min et al., 2022). This process ensures that the data-specific weights $w^{u}_{k}$ reflect interpretable insights from the literature on graph properties, GNNs, and performance, aligning with general empirical evidence to construct trustworthy meta-knowledge that empowers LLMs to quickly design effective, data-aware models for unseen tasks.
@@
Knowledge Pool Construction. Using the computed task similarity set $\mathcal{S}^{u}$ for $G^{u}$, we rank the benchmark graphs and select the Top-$N_{s}$ most similar sources. The knowledge pool $\mathcal{K}$ for $G^{u}$ is then constructed by aggregating the top-performing models from them:
(6) $\mathcal{K}=\bigcup_{G^{i}\in\text{Top-}{N_{s}}(\mathcal{S}^{u})}\{(G^{i},\{\theta_{im}^{*}\}_{m=1}^{N_{m}})\},$
@@
The top-performing models $\{\theta_{i}^{*}\}$ act as a truncated prior $\mathbb{P}(\theta_{i}^{*})$, summarizing accumulated design knowledge from benchmark $G^{i}$, while task similarity $\mathcal{S}(G^{u},G^{i})$ serves as a gating module for solving pseudo-likelihood $\mathbb{P}(\mathcal{F}(G^{u})|\theta_{i}^{*})$, indicating the effectiveness of $\theta_{i}^{*}$ on unseen graph $G^{u}$.
@@
Table 1. Initial performance comparison (accuracy) of GNNs designed by different methods. $N_{s}=3$ for Desi-Init. Best overall results are in bold, and best among each category are underlined.
@@
3.3. Model Suggestion and Refinement
Overall, our formulation of Knowledge Pool $\mathcal{K}$ approximates the posterior $\mathbb{P}(\theta_{i}^{*}|\mathcal{F}(G^{u}))$, prioritizing models that align with data-aware evidence. This Bayesian-inspired process ensures solid use of constructed meta-knowledge for designing models in both one-shot and a few attempts.
Initial Model Suggestion. By supplying contextual meta-knowledge of high similarities to the unseen graph $G^{u}$ from $\mathcal{K}$, we leverage LLM as a surrogate meta-controller to sample $N_{s}$ initial proposals $\{\theta_{ui}\}_{i=1}^{N_{s}}$ based on the approximated posterior, where $\mathcal{K}_{i}$ represents the specific knowledge from the $i$-th similar benchmark graph in $\mathcal{K}$: $\theta_{ui}\leftarrow\mathcal{LLM}_{IMS}(\mathcal{F}(G^{u},\bar{I}),\mathcal{K}_{i}).$. Compared to existing AutoGNNs, our method avoids any model training on unseen data, delivering promising designs by combining fine-grained literature and empirical insights for immediate model deployment.
Knowledge-Driven Refinement. To further optimize initial model proposals, we design a structured, knowledge-driven refinement strategy inspired by expert-like exploration-exploitation trade-offs, as illustrated in Figure 2. Our local search strategy iteratively improves models using the constructed meta-knowledge of high similarities to the unseen from $\mathcal{K}$. (0) Re-Ranking: Initial proposals $\{\theta_{ui}\}_{i=1}^{N_{s}}$ are re-ranked based on their validation performance, and the best-performing proposal $\theta_{u1}$ serves as the refinement starting point. (1) Controlled Exploration: Configurations from other models in $\mathcal{K}$ are crossovered with $\theta_{u1}$ to generate $N_{c}$ new candidates. (2) Model Promotion Mechanism: Candidates are ranked based on their retrieved performances on the benchmark dataset of $\mathcal{K}_{1}$, with the most promising candidate $\theta_{u}^{t^{\prime}}$ advanced for further refinement at iteration $t$. (3) Directional Exploitation: The LLM meta-controller $\mathcal{LLM}_{KDR}$ mutates $\theta_{u}^{t^{\prime}}$ using user requirements, task descriptions, search space details, and previous training logs. (4) Evaluation and Update: Each refined candidate $\theta_{u}^{t}$ is validated and added to the optimization trajectory $\Theta^{T}$. The best-performing model is updated as $\theta_{u}^{*}$. This refinement process balances exploration (testing diverse configurations) and exploitation (focusing on high-potential candidates), leveraging prior knowledge in the textual space to efficiently optimize model designs.
4. Experiments
4.1. Experimental Settings
Task and Datasets. We conduct node classification studies using 8 benchmark datasets from NAS-Bench-Graph (Qin et al., 2022): Cora, Citeseer, PubMed (Sen et al., 2008), CS, Physics, Photo, Computer (Shchur et al., 2018), and ogbn-arXiv (Hu et al., 2020). We include 3 extra datasets for out-of-distribution (e.g., heterophily) generalization: DBLP (Bojchevski and Günnemann, 2017), Flickr (Zeng et al., 2019), and Actor (Pei et al., 2020).
Baselines. We compare DesiGNN with 4 categories of 17 baselines: (1) Manually Designed GNNs: 7 GNNs (Kipf and Welling, 2016; Veličković et al., 2017; Hamilton et al., 2017; Xu et al., 2018; Defferrard et al., 2016; Bianchi et al., 2021; Morris et al., 2019) showcasing human experts’ efforts in designing tailored models for specific types of graphs; (2) Classical Automated Approaches: 3 typical automated search strategies (Li and Talwalkar, 2020; Real et al., 2019; Zoph and Le, 2016), plus 2 advanced AutoGNN systems (Gao et al., 2021; Zhou et al., 2022), all operating within the same search space as DesiGNN for fairness; (3) Meta-based Approaches: 3 meta-based, semi-supervised methods (You et al., 2020; Qin et al., 2022; Cao et al., 2023) that recommend models from the most similar benchmark graphs or based on meta-knowledge after testing anchor models on the unseen task; (4) LLM-Based Approaches: 2 recent SOTA LLM-Based methods (Wang et al., 2023a; Dong et al., 2023) that leverage LLMs for iterative architecture search.
Evaluation Metrics. To ensure reliability, we report the average accuracy and standard deviation over 10 runs. For automated methods, we also analyze the best-so-far accuracy after validating 1-30 model proposals to measure short-run efficiency. To assess efficiency under varying LLM API traffic, we set the basic time unit as the time to validate a model proposal on the unseen dataset.
Implementation Details. Our code [^1] Full code, data, prompts: https://github.com/jilwang84/DesiGNN. is implemented in PyTorch and LangChain (GPT-4). For clarity, Desi-Init refers to initial model proposals, while DesiGNN denotes the whole pipeline, including model refinement. We use NAS-Bench-Graph (Qin et al., 2022) to accumulate knowledge and provide a unified search space. To prevent data leakage, when a benchmark dataset is treated as unseen, it is completely excluded from the knowledge retrieval process and anonymized to ensure fair evaluation. We have validated that LLMs, given only dataset descriptions, lack prior knowledge of effective architectures in the NAS-Bench-Graph search space.
Figure 5. Short-run performance of DesiGNN compared to stronger automated baselines.
4.2. Main Results